OpenStudy (anonymous):

Medal A physician orders 20mcg teriparatideSCq.d. The stock vial contains 750 mcg /15mLsolution. Interpret the order and determine how much of the solution to administer.

OpenStudy (anonymous):

50mcg? im not sure

OpenStudy (anonymous):

how did you get this?

OpenStudy (anonymous):

@incognlto do you understand this

OpenStudy (lncognlto):

I don't ;-;

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@Lurker do you understand this answer?

OpenStudy (lurker):

okay so again we need the same ratios

OpenStudy (lurker):

750 mcg/15mL we only need 20mcg how many mL will that be in 20/x=750/15

OpenStudy (anonymous):

so 20/x=50

OpenStudy (anonymous):

300/750

OpenStudy (anonymous):

thats 0.4

OpenStudy (anonymous):

yes 0.4mcg

OpenStudy (anonymous):

is that right? @Lurker

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

medals?

OpenStudy (anonymous):

is in the answer in mL or mcg?